OpenStudy (anonymous):

Where can I find the formulas for calculating the results of two balls colliding where both are moving? The collision is elastic.

OpenStudy (anonymous):

Well what are you trying to calculate? The velocity before and after the collision? The energy of the balls before and during the collision? You have to be more specific as to what kind of calculations you are looking for.

OpenStudy (anonymous):

Two billiard balls roll towards each other. They each have a mass of 0.3Kg. Ball one is moving at V1=1m/s to the right, while ball 2 is moving at v2=0.8m/s to the left. Analyse the system and find its total momentum and the new direction of travel.

OpenStudy (anonymous):

That is the exact question I need to answer.

OpenStudy (anonymous):

Well you know that momentum (p) is equal to mass multiplied by velocity. So we have the equation p=m|v| where m is equal to the total mass of the system (0.3+0.3=0.6Kg) and v is equal to the total net velocity of the system (0.1+(-0.8)=-0.7. All the negative sign indicates in this case is that the system is moving in the direction of the ball with velocity v2 so it's moving to the left. When we calculate for momentum we get p=m|v|=(0.6Kg)*|(-0.7m/s)| = 0.42Kgm/s

OpenStudy (anonymous):

Thanks so much mate, Your a life saver. :)

OpenStudy (anonymous):

The unit is Kilogram-Meters per Second. Just in case you didn't know. :P